OpenStudy (anonymous):

find an equation of tangent to the graph of f(x)= x^3-2x at x=1

OpenStudy (goformit100):

Fin the tangent of the function

OpenStudy (goformit100):

BY differentiating the function you can find the tangent

OpenStudy (anonymous):

Suppose y= x^3-2x differentiating it, dy/dx= 3x^2-2x substituting x=1 which gives us dy/dx= 1 substituting x=1 in the given curve equation to get y which is equal to -1 Use the formula 'y-y1=dy/dx(x-x1) Where y1=-1 and x1=1 equation of tangent, y=x-2 :)